Suppose that the mean number of students who call me during by




sleepy hours ( 2pm-3pm) is 5.




What is the probability that in a given hour, exactly two students will call?




And what is the probability that more than two will call in the given hour?

1
Expert's answer
2021-12-12T18:24:26-0500

This is a Poisson distribution with "\mu=5."

"P(X=2)=e^{-5}\frac{5^2}{2!}=0.0842."

"P(X>2)=1-P(X\le2)=1-P(X=0)-P(X=1)-P(X=2)="

"=1-e^{-5}(\frac{5^0}{0!}+\frac{5^1}{1!}+\frac{5^2}{2!})=0.8754."